Math Question says Estimate the solution for the proportions. Example m/25=16/98
Answer by stanbon(48568)

Math Question says Estimate the solution for the proportions. Example m/25=16/98
---------------------
98 is about 6 times 16
---------
So m should be about 1/6 th of 25 or around 4
-------------------
What is the exact answer?
======================
m = 25(16/98)
m = 400/98
m = 4.08..
Cheers,
Stan H.